Suppose the demand curve is D(p)=100-2p. What price would the monopolist set if he had 60 apartments? How many would he rent? What price would he set if he had 40 apartments? How many would he rent? (Assume that the monopolist faces no costs)
Answer #1

Monopoly Qd=100-2P

P=50-0.5Qd

TR=P*Q=50Q-0.5Q2

Total revenue is maximised when MR=dTR/dQ=50-Q

Thus Profit is maximised when Q=50 and price=25

1) Thus monopoly will charge price=25 and sells only 50 apartments and profit=1250

2) When apartments are only 40 in number then Demand=100-2P=40

Price=(100-40)/2=30 and profit=30*40=1200